The perimeter of a Desktop is 108 inches, the width is 14 inches shorter than the length,write an eqaution to represent the perimeter and determine the dimension of the desktop?
Answer by checkley77(12844)

W=L-14
2L+2W=108
2L+2(L-14)=108
2L+2L-28=108
4L=108+28
4L=136
L=136/4
L=34 ANS. FOR THE LENGTH.
W=34-14=20 ANS. FOR THE WIDTH.
PROOF:
2*34+2*20=108
68+40=108
108=108